Matematica, Anno Accademico 2006-2007, Scienze Geologiche M. Novaga, V.M. Tortorelli I prova in itinere 6 ottobre 2006 RISULTATI

ALDOVARDI Gianfranco 6/7 ...... n BALDI Lorenzo 6/7......vale BALLONI Francesca 6/7......n BANCI Federico 3/4...... BARANDONI Andrea 6/7 ......n BARONDI Simone 6/7......vale BARRUFFI Anna Chiara 6/7......n BARTOLI Fabrizio 6/7......suff BEDINI Federica ?......n BENDINELII Filippo 6/7......n BIAGIONI Riccardo 6/7......*suff BLASI Matteo 6/7...... BONAVENTURI David 6/7......n BORGHINI Davide 3/4...... BRONDANI Andrea 3/4...... CAMPENNI Alessio 6/7 ...... n CARMIGNANI GuglilmoTommaso 6/7 ......n CASACCHIA Claudia 5/6......n CAVATI Cristiano 6/7......n CECCARELLI Dario 4/5......suff CHINI Alessia 6/7......vale CIANI Valentina 4/5......suff+ CIGNONI Giulia 6/7......n CINQUINI Ilaria 6/7......vale CIUFI Roberta 3/4...... COLUCCINI Leonardo 6/7......n DA MOMMIO Alessandro 4/5? ......*vale DA PRATO Giulia 6/7......vale DI IORIO Jacopo 6/7......n FAZIO Agnese 6/7......vale FRANCESCHI Clara 6/7......suff GAMBINO Mario 6/7......vale GENOVESI Matteo 6/7......n GISFREDI Chiara 6/7......vale GORI Francesco ?6/7...... GUARGUAGLINI Matteo 6/7......n LOIACONI Rocco 3/4...... MADRIGALI Paolo 6/7......n MARCHI Alessandra 6/7......*vale MARCHINI Davide 6/7...... MARIOTTI Sara 6/7......n MEINI Luca 6/7......n MICHELI Giovanni 6/7......n MILAZZO Flavio 6/7.......n NENCI Alessandro 6/7......vale-NIRTA Antonio 5/6......n NISTICO’ Rosaria 2/3 ...... OBEGI MONZER 6/7......vale PACINI Dario 4/5......*n+ PANCANI Dario 6/7......n PARDINI Dario 4/5......*suff PARLANTI Benedetta 5/6...... PENNINO Matteo 6/7......n PENTANGELO Paolo 6/7......vale-PIERUCCIONI Diego 6/7......n PINELLI Gabriele 6/7......vale PUCCI Luca 5/6...... RALLO Maurizio 6/7......vale REDINI Federica 6/7...... RIELLI Andrea 6/7......n RIZZO Serena 4/5...... ROCCHICCIOLI Samuel 5/6......vale ROTONDI Gabriele ?......suff-SCATENI Bianca 4/5...... TAGLIAMONTE Cesare 6/7......n TESCONI Simone 6/7......n TINE Alice 6/7......n VANNUCCI Marcella 6/7......n VANNUCCI Simone ?.....*.n VITI Federico 4/5......*n VIVIANI Diego 6/7......n